# Effect of Domain Generalization Techniques in Low-Resource Systems

### v1

@@ 3.2 Invariance via Causal Representation Learning
Invariance via Causal Representation Learning (CRL) during inference is one of the causal domain generalization techniques we aim to explore in low-resource settings, as outlined by Liu et al. [1]. Our goal is to examine how domain generalization (DG) can be effectively applied under such conditions. To achieve this, we adopt the DINER: Debiasing Aspect-based Sentiment Analysis with Multi-variable Causal Inference[4], a debiasing approach for the Aspect-Based Sentiment Analysis (ABSA) task. We further discuss DINER and our methodology in subsequent sections.
Figure 1 demonstrates how high-capacity neural networks can easily learn spurious correlations arising from annotation bias. These biases cause models to capture misleading dependencies rather than true causal relationships. Due to their high representational power, neural models are particularly vulnerable to such biases, often learning shortcuts instead of underlying latent features [21, 22, 23]. For instance, reversing the polarity of target aspects can result in up to a 20% drop in accuracy, highlighting the model’s reliance on these correlations. The figure also supports the observation that models frequently depend on spurious correlations between target and non-target aspects in sentiment datasets such as Pontiki et al. [5]. In the ’Original’ setting, both review-only and aspect-only models perform well; however, when these correlations are disrupted in the ’RevTgt’, ’RevNon’ and ’AddDiff’ settings, performance decreases or becomes balanced. This confirms that neural models often exploit dataset biases rather than learning genuine target-specific sentiment representations.
3.2.1 Debiasing in Aspect Review using DINER
To mitigate these effects, recent studies have proposed several debiasing strategies, including argumentation-based methods[19], reweighting-based training[24], and causal inference-based approaches[25, 1]. Among these, causal inference has attracted significant interest for its solid theoretical foundation and minimal disruption to existing learning paradigms.
The DINER framework[4] models ABSA as a Structural Causal Model (SCM) with five key variables: confounder $C$ (prior contextual knowledge), review $R$, aspect $A$, fused knowledge $K$ (derived from $R$ and $A$), and sentiment label $L$. The causal relations are defined as $R\rightarrow K\leftarrow A$, $K\rightarrow L$, $R\rightarrow L\leftarrow A$, and $C\rightarrow R,C\rightarrow L$. The causal structure is illustrated in Figure 2, showing the directed acyclic graph (DAG) that represents relationships among these variables. The objective of DINER is to isolate the causal path $K\rightarrow L$ while eliminating spurious correlations via $R\rightarrow L\leftarrow A$ and mitigating confounding through $C\rightarrow R$ and $C\rightarrow L$. For details on causal graph construction, please refer to Wu et al. [4].
Figure 2: Causal structure of the DINER framework ([4]).
The debiased review representation is computed as:
[…]
$\begin{split}TIE_{a,r}&=TE_{a,r}-NDE_{r}-NDE_{a}+IE_{a,r}\\ &=TE_{a,r}-(NDE_{r}+NDE_{a})\quad(\text{since }IE_{a,r}=0)\end{split}$ (7)
Following Niu et al. [25], we define:
Thus:
The final prediction is implemented through a model ensemble:
@@ 3.2.2 Data
We used Afri-SemEval, a multilingual dataset created by translating the SemEval 2014 [5]) dataset into 17 African languages. The dataset spans diverse languages from multiple African language families, as shown in Table 1, [29].
Each translated instance preserves the original structure of SemEval, including the review text, target aspect term, and sentiment polarity (posistive, negative, neutral). Translation challenges included maintaining aspect-term alignment, preserving sentiment polarity, and capturing cultural nuances in sentiment expression.
Table 1: List of 17 African supported languages
@@ 9 Kirundi run
3.3 African Language Adaptation
A major challenge in adapting DINER to African languages lies in aspect term alignment. The original framework assumes a direct mapping between aspect terms and their positions in the review, but translation often disrupts this due to word order variation (e.g., English SVO vs. Hausa SXV), morphological complexity (e.g., noun classes in Bantu languages), and differences in expression length (multiword to single-word mappings, and vice versa).
We address this with an alignment heuristic based on semantic similarity. Embeddings are computed for translated aspect terms and all n-grams in the review. The n-gram with the highest similarity to the aspect term is then selected as its aligned position for DINER processing. While effective for single-token aspects, this heuristic remains challenging for multi-token and non-contiguous aspect terms, which are common in many African languages.
@@ 4 Experiment and Result
4.1 Invariance via Causal Data Augmentation Results
@@
where $h_{\theta}$ denotes the trained classifier. A smaller cross-domain difference indicates stronger invariance to domain shift induced by paraphrastic augmentation.
@@ 4.1.1 Results on Yoruba
Tables 2 and 3 summarize the validation and test accuracies for Yoruba, respectively. The combined configuration consistently yields higher accuracy across domains, confirming that integrating both original and paraphrased data enhances generalization.
Table 2: Validation results on Yoruba subset (Logistic Classifier).
@@
Table 3: Test results on Yoruba subset (Logistic Classifier).
@@ 4.1.2 Results on Igbo
Tables 4 and 5 present the corresponding results for the Igbo subset. The general tend in this case, does not mirror Yoruba. While the combined dataset achieves stable performance across domains, improvements are more pronounced on our original data due to a larger semantic drift between original and paraphrased data.
Table 4: Validation results on Igbo subset (Logistic Classifier).
@@
Table 5: Test results on Igbo subset (Logistic Classifier).
@@
Overall, results confirm that causal-preserving data augmentation yields measurable improvements in cross-domain robustness. In Yoruba, where paraphrases were semantically well-aligned, the combined model achieved the best generalization with the smallest invariance gap $\Delta_{\text{inv}}$. In Igbo, where GPT-4o paraphrases introduced a larger distributional shift, the combined model still maintained competitive performance, albeit with smaller gains.
These findings support the hypothesis that training on both original and paraphrased data enhances model invariance to superficial linguistic variability. The modest but consistent improvements, particularly in Yoruba, shows the potential of causal data augmentation as a lightweight strategy for generalization in low-resource sentiment classification.
@@ 4.2 Invariance via Causal Representation Learning
4.2.1 Training Configuration
We evaluate three transformer-based models: XLM-R, a multilingual RoBERTa model trained on 100 languages[30] ; Afro-XLMR Large, a specialized version fine-tuned for African languages[26] and Afro-XLMR Large-76L[31], an extended variant with additional African language coverage. For each model, we compare two training configurations: CT=False, which implements standard training where aspect and review are concatenated, and CT=True, which employs causal training where aspect and review are treated separately.
All models were fine-tuned with consistent hyperparameters including a learning rate of 2e-5, batch size of 16, maximum sequence length of 128, 10 training epochs, weight decay of 0.01, and the AdamW optimizer. For CT=True models, we implemented DINER’s causal inference components including backdoor adjustment with $\tau=0.1$, counterfactual reasoning using shared encoders, and TIE calculation with ensemble fusion function.
We evaluate performance using accuracy, macro F1-score, and Aspect Robustness Score (ARS), which measures performance under aspect perturbations[4]. To assess out-of-domain (OOD) generalization, we employ cross-language evaluation by training on one language and testing on others, as well as perturbation tests including RevTgt (reversing the polarity of target aspects), RevNon (reversing the polarity of non-target aspects), and AddDiff (adding aspects with different polarity).
4.2.2 Results on Convergence Analysis
A key finding is that causal training (CT=True) typically converges faster across multiple transformer models. As shown in our accuracy charts (Figures 3(a)-3(c)), models trained with causal inference require fewer steps to reach comparable performance. XLM-R with CT=True converges 13 steps earlier than CT=False, while Afro-XLMR Large with CT=True shows steeper initial learning curves, and Afro-XLMR Large-76L exhibits similar patterns with accelerated convergence. This efficiency advantage is particularly valuable in resource-constrained African AI contexts, enabling faster training cycles and reduced computational demands.
(a) XLM-R accuracy
(b) Afro-XLMR-large accuracy
(c) Afro-XLMR-large-76L accuracy
Figure 3: Comparison of accuracy charts. (a) XLM-R, (b) Afro-XLMR-large, (c) Afro-XLMR-large-76L.
4.2.3 Results on Performance Across Languages
Performance varies significantly across African languages, reflecting both linguistic differences and translation quality. Table 6 shows that for the 5 low-resource languages studied in depth, results differ considerably. Afrikaans demonstrates the highest overall performance, likely due to its similarity to English and high-quality translations. Amharic shows the lowest performance across all models, reflecting script and morphological complexity. Hausa exhibits moderate performance with significant gains under causal training. Igbo shows notable improvement with CT=True in XLM-R but less pronounced gains in AfroXLM variants. Yoruba follows a similar pattern to Igbo, with causal training providing modest improvements. These results highlight the language-dependent nature of causal inference benefits, with some languages benefiting more than others.
@@ 4.2.3 Results on Performance Across Languages
4.2.4 OOD Generalization Results
The theoretical advantage of causal models in OOD settings shows mixed results in practice. Table 7 reveals that XLM-R shows better generalization with CT=True for Afrikaans, Igbo, and Yoruba, while AfroXLM-R Large underperforms with CT=True on many OOD tasks, though exceptions exist where causal training significantly improves OOD performance. These inconsistencies likely stem from translation noise affecting causal variable alignment, model architecture sensitivity to causal training, and varying degrees of distributional shift across languages.
Table 7: Performance in OOD settings (Accuracy %)
@@ 4.2.4 OOD Generalization Results
4.2.5 Case Studies
Table 8 presents illustrative examples demonstrating the robustness improvements from causal training. In the original example, both baseline and DINER correctly classify positive sentiment. For RevTgt cases, both models correctly adjust when target sentiment is reversed. However, for RevNon scenarios, DINER correctly maintains focus on target aspect despite negative non-target aspects, while baseline models fail. Similarly, with AddDiff examples, DINER shows robustness when additional negative aspects are introduced. These examples highlight DINER’s ability to isolate target aspect sentiment from confounding context—a crucial capability for practical ABSA applications.
Table 8: Example case studies. The corresponding gold labels and the predictions for each example are presented.
@@ 4.2.5 Case Studies
AddDiff The food is top notch, the service is attentive, and the atmosphere is great, but music is too heavy, waiters is angry and staff is arrogant. Positive Negative ✗ Positive ✓
5 Conclusion and Future Work
In this study, we systematically investigated the efficacy of two distinct causal domain generalization techniques in the challenging context of low-resource African languages. By examining both a data-level intervention through Causal Data Augmentation (CDA) and a representation-level intervention with Invariant Causal Representation Learning (ICRL), we aimed to understand their complementary strengths and practical trade-offs.
Our findings confirm that both approaches offer tangible benefits for enhancing model robustness against distribution shifts. The CDA experiments on the NaijaSenti corpus demonstrated that augmenting training data with semantically equivalent paraphrases is a lightweight yet effective strategy for improving cross-domain generalization. The positive results in Yoruba, where the combined model achieved higher accuracy and a smaller invariance gap, underscore the potential of this method when paraphrase quality is high. Conversely, the results in Igbo highlight a key dependency: the success of CDA is contingent on the semantic fidelity of the generated counterfactuals, as significant distributional drift can temper performance gains.
The application of the DINER framework to our newly created Afri-SemEval benchmark provided a more nuanced picture. While causal training consistently accelerated model convergence and improved in-domain accuracy for many of the 17 languages, its impact on Out-of-Distribution (OOD) generalization was varied. This inconsistency reveals a critical insight: the theoretical promise of sophisticated causal models can be attenuated by the practical realities of cross-lingual transfer. We identify a "dual-shift" problem, where models must contend not only with the intended linguistic domain shift but also with an unintended, artifact-based shift introduced by translation noise and aspect alignment errors. This suggests that the language-dependent nature of the observed gains is a direct reflection of this complex interaction.
Looking ahead, our work points to several promising research directions. First, the limitations of our heuristic-based aspect alignment highlight the need for more robust methods to adapt complex models like DINER to typologically diverse languages. Future work could explore jointly trained alignment models or the development of syntactic parsers tailored to African languages. Second, the variable quality of LLM-generated paraphrases suggests a need for more controlled generation techniques or human-in-the-loop validation pipelines to maximize the benefits of CDA. Finally, the "dual-shift" phenomenon calls for the development of novel domain generalization methods that are explicitly designed to be robust not only to natural linguistic variation but also to the systematic noise inherent in cross-lingual transfer. A hybrid approach, perhaps using CDA to create more robust pre-trained models before applying representation-level debiasing, could prove to be a fruitful avenue for future exploration.
Acknowledgments:
We are grateful to ML Collective for generously providing the computational resources required for this study. The success of our extensive experiments, particularly the resource-intensive large language model (LLM) fine-tuning and the multi-language DINER training, was directly enabled by their dedicated compute support.

### v2

@@ 3.2 Invariance via Causal Representation Learning
Invariance via Causal Representation Learning (CRL) during inference is one of the causal domain generalization techniques we explore in low-resource settings. To achieve this, we adopt the DINER: Debiasing Aspect-based Sentiment Analysis with Multi-variable Causal Inference [17], a debiasing approach for the Aspect-Based Sentiment Analysis (ABSA) task. We discuss our methodology in what follows.
3.2.1 Debiasing in Aspect Review (DINER)
Recent studies have proposed several debiasing strategies, including argumentation-based methods [20], reweighting-based training [21], and causal inference-based approaches [22, 23]. Among these, causal inference has attracted significant interest for its solid theoretical foundation and minimal disruption to existing learning paradigms. See Section A for more details on debiasing in aspect review using DINER.
@@ 3.2.2 Data
We introduce Afri-SemEval, a multilingual dataset created by translating the SemEval 2014 dataset into 17 African languages to support research on domain generalization in low-resource contexts. Figure 4 illustrates our dataset creation pipeline. We leverage the Microsoft Azure Translator API for one-to-many translation, where English reviews are batch-processed with configurations settings to comply with API constraints and translated simultaneously into all target languages. The pipeline parses JSON responses, maps translations to source review IDs, and applies a semantic similarity heuristic to align aspect terms across languages. The dataset spans diverse languages from multiple African language families, as shown in Table 3.
Figure 4: Afri-SemEval dataset creation pipeline showing the translation workflow for SemEval-2014 to 17 African languages using Azure API.
@@ 9 Kirundi run
A major challenge in adapting DINER to African languages lies in aspect term alignment. The original framework assumes a direct mapping between aspect terms and their positions in the review, but translation often disrupts this due to word order variation, morphological complexity (e.g., noun classes in Bantu languages), and differences in expression length (multiword to single-word mappings, and vice versa).
We address this with an alignment heuristic based on semantic similarity. Embeddings are computed for translated aspect terms and all n-grams in the review. The n-gram with the highest similarity to the aspect term is then selected as its aligned position for DINER processing. While effective for single-token aspects, this heuristic remains challenging for multi-token and non-contiguous aspect terms.
@@ 4 Experiment and Result
4.1 Invariance via Causal Data Augmentation
@@
where $h_{\theta}$ denotes the trained classifier.
@@ 4.1.1 Results on Yoruba
Tables 4 and 5 summarize the validation and test accuracies for Yoruba, respectively. The combined configuration consistently yields higher accuracy across domains, confirming that integrating both original and paraphrased data enhances generalization.
@@
Table 4: Validation results on Yoruba subset.
@@ 4.1.2 Results on Igbo
Tables 6 and 7 present the corresponding results for the Igbo subset. The general trend in this case, does not mirror Yoruba. While the combined dataset achieves stable performance across domains, improvements are more pronounced on our original data due to a larger semantic drift between original and paraphrased data.
@@
Table 6: Validation results on Igbo subset.
@@
Overall, our results confirm that causal-preserving data augmentation yields measurable improvements in cross-domain robustness. In Yoruba, where paraphrases were semantically well-aligned, the combined model achieved the best generalization with the smallest invariance gap $\Delta_{\text{inv}}$. In Igbo, where paraphrases had a larger distributional shift, the combined model still maintained competitive performance, with smaller gains. These findings suggests that training on both original and paraphrased data enhances model invariance to superficial linguistic variability. The modest but consistent improvements show the potential of causal data augmentation as a lightweight strategy for generalization in low-resource sentiment classification.
@@ 4.2 Invariance via Causal Representation Learning
For our experiments, we evaluated three transformer-based models: XLM-R [24], Afro-XLMR Large [25], and Afro-XLMR Large-76L [26]. For each model, we compare two training configurations: CT=False (standard training with concatenated aspect and review) and CT=True (causal training with separate aspect and review processing).
All models were fine-tuned with learning rate 2e-5, batch size 16, maximum sequence length 128, 10 epochs, weight decay 0.01, and AdamW optimizer. For CT=True models, we implemented DINER’s causal inference components including backdoor adjustment ($\tau=0.1$), counterfactual reasoning, and TIE calculation with ensemble fusion.
We use accuracy as metrics to access the performance of our approach. To assess out-of-distribution (OOD) generalization, we employ cross-language evaluation and perturbation tests including RevTgt (reversing target aspect polarity), RevNon (reversing non-target aspect polarity), and AddDiff (adding aspects with different polarity).
4.2.1 Results
A key finding is that causal training (CT=True) typically converges faster across multiple transformer models. As shown in Figures 5, 6, and 7, models trained with causal inference require fewer steps to reach comparable performance. This efficiency advantage is particularly valuable in resource-constrained contexts, enabling faster training cycles and reduced computational demands.
Figure 5: XLM-R accuracy comparison showing convergence patterns for CT=False and CT=True configurations.
Figure 6: Afro-XLMR-large accuracy comparison showing convergence patterns for CT=False and CT=True configurations.
Figure 7: Afro-XLMR-large-76L accuracy comparison showing convergence patterns for CT=False and CT=True configurations.
Performance varies across African languages, reflecting both linguistic differences and translation quality. Table 8 presents the performance results across five low-resource African languages for three transformer-based models under both standard training (CT=False) and causal training (CT=True) configurations. Afrikaans demonstrates the highest overall performance, likely due to its similarity to English. Amharic shows the lowest performance, reflecting script and morphological complexity. Hausa, Igbo, and Yoruba exhibit moderate performance with varying gains under causal training.
@@ 4.2.1 Results
The theoretical advantage of causal models in OOD settings shows mixed results in practice. Table 9 shows the OOD generalization results across the same five languages, demonstrating varying effectiveness of causal training across different model architectures. The results reveals that XLM-R generally benefits from CT=True for Afrikaans, Igbo, and Yoruba, while AfroXLM-R Large shows inconsistent patterns. These inconsistencies likely stem from translation noise affecting causal variable alignment and varying degrees of distributional shift across languages.
@@
Table 9: Performance in OOD settings (Accuracy %)
In section  A.3, we further discussed some examples and case studies that demonstrates the robustness improvements from the causal training in our CRL experiments. See table  10.
5 Conclusion and Future Work
In this study, we investigated two causal domain generalization techniques for low-resource African languages: Causal Data Augmentation (CDA) and Invariant Causal Representation Learning (ICRL) via the DINER framework.
Our findings confirm that both approaches enhance model robustness against distribution shifts. CDA experiments on NaijaSenti showed that augmenting with semantically equivalent paraphrases improves cross-domain generalization, particularly in the Yoruba dataset where paraphrase quality was high. In contrast, Igbo results highlight that success depends on semantic fidelity, as distributional drift can limit gains. Applying DINER to the Afri-SemEval dataset demonstrated that causal training accelerates convergence and improves in-domain accuracy across 17 languages, while out-of-distribution generalization varied. This variability reveals a dual-shift problem, where models face both linguistic domain shifts and artifacts introduced by translation noise and aspect misalignment.
Future work should address three directions: developing robust aspect alignment methods for typologically diverse languages, improving LLM-generated paraphrase quality through controlled generation or human-in-the-loop validation, and designing domain generalization methods robust to both natural linguistic variation and cross-lingual transfer noise. A hybrid approach combining CDA-enhanced pretraining with representation-level debiasing may be effective.
Ethical Considerations This work focuses on low-resource languages, which are historically underrepresented in NLP. While our dataset and models aim to improve language coverage and model generalization in distribution shift settings, they may still reflect cultural or social biases present in the source data.
We only use publicly available text and synthetic paraphrases, avoiding private or sensitive information. Finally, generated paraphrases should respect cultural norms to avoid offensive or misleading outputs, ensuring responsible and ethical use of the resources we provide.
Limitations While this study demonstrates the effectiveness of causal domain generalization techniques in low-resource NLP, several limitations remain. First, the Afri-SemEval dataset relies on machine translation, which may introduce artifacts and fail to capture nuances of native African language expressions. Second, our experiments focus on sentiment analysis, and the generalizability of these findings to other NLP tasks remains unexamined. Third, the computational cost of counterfactual data augmentation using large language models may limit scalability in resource-constrained settings. Finally, our evaluation covers only 17 African languages, and broader linguistic coverage would strengthen the conclusions.
Acknowledgment:
We are grateful to ML Collective for generously providing the computational resources required for this study. The success of our extensive experiments, in particular, the resource-intensive large language model (LLM) fine-tuning and the multi-language DINER training, was directly enabled by their dedicated compute support.
Appendix A Appendix
A.1 DINER FRAMEWORK
The DINER framework models ABSA as a Structural Causal Model (SCM) with five key variables: confounder $C$ (prior contextual knowledge), review $R$, aspect $A$, fused knowledge $K$ (derived from $R$ and $A$), and sentiment label $L$. The causal relations are defined as $R\rightarrow K\leftarrow A$, $K\rightarrow L$, $R\rightarrow L\leftarrow A$, and $C\rightarrow R,C\rightarrow L$. The causal structure is illustrated in Figure 9, showing the directed acyclic graph (DAG) that represents the relationships between these variables. The objective of DINER is to isolate the causal path $K\rightarrow L$ while eliminating spurious correlations through $R\rightarrow L\leftarrow A$ and minimizing confounding through $C\rightarrow R$ and $C\rightarrow L$. For details on causal graph construction, see [17].
A.2 Causal Variables Discovery
Figure 8 demonstrates how high-capacity neural networks can easily learn spurious correlations arising from annotation bias. These biases cause models to capture misleading dependencies rather than true causal relationships. Due to their high representational power, neural models are particularly vulnerable to such biases, often learning shortcuts instead of underlying latent features ([27, 28, 29]). For instance, reversing the polarity of target aspects can result in up to a 20% drop in accuracy, highlighting the model’s reliance on these correlations. The figure also supports the observation that models frequently depend on spurious correlations between target and non-target aspects in sentiment datasets such as [8]. In the “Original” setting, both review-only and aspect-only models perform well; however, when these correlations are disrupted in the “RevTgt,” “RevNon,” and “AddDiff” settings, performance decreases or becomes balanced. This confirms that neural models often exploit dataset biases rather than learning genuine target-specific sentiment representations.
Figure 9: Causal structure of the DINER framework ([17]).
@@ A.2 Causal Variables Discovery
AddDiff The food is top notch, the service is attentive, and the atmosphere is great, but the music is too heavy and the staff is arrogant. Positive Negative ✗ Positive ✓
Table 10: Illustrative case studies showing model robustness under target and non-target sentiment perturbations.
The debiased review representation is computed as:
[…]
$\begin{split}TIE_{a,r}&=TE_{a,r}-NDE_{r}-NDE_{a}+IE_{a,r}\\ &=TE_{a,r}-(NDE_{r}+NDE_{a})\\ &\quad(\text{since }IE_{a,r}=0)\end{split}$ (6)
Following [22], we define:
Thus:
The final prediction is implemented through a model ensemble ([17]):
A.3 Case Studies
Table 10 presents illustrative examples demonstrating the robustness improvements from our causal training. In the original example, both baseline and DINER correctly classify positive sentiment. For RevTgt cases, both models correctly adjust when target sentiment is reversed. However, for RevNon scenarios, DINER correctly maintains focus on target aspect despite negative non-target aspects, while baseline models fail. Similarly, with AddDiff examples, DINER shows robustness when additional negative aspects are introduced. These examples highlight DINER’s ability to isolate target aspect sentiment from confounding context—a crucial capability for practical ABSA applications.